How do you find what w is in this problem :
39-12w=7-16w

Answer by edjones(7569)

39-12w=7-16w
get the w's on one side and the rest on the other by performing an operation on both sides.
39-12w+16w=7-16w+16w
4w+39-39=7-39
4w=-32
4w/4=-32/4
w=-8
Ed